OpenStudy (anonymous):

The two square pyramids are similar. Find the total volume of both pyramids if the ratio of their surface areas is 9/16

OpenStudy (anonymous):

SINCE THE TWO PYRAMIDS ARE SIMILAR THE RATIO OF THEIE SURFACE AREAS =9/16 HENCE THE RATIO OF THE EDGES WILL BE SQRT(9/16) =3/4 AND THE RATIO OF THEIR VOLUMES WILL BE (3/4)^3 =27/64 HOPE NOW YOU CAN DO..

OpenStudy (imstuck):

Area is a squared measurement, so that means that if their areas are 9 and 16, those are squared numbers. In order to find their volumes, you have to work backwards to find their similarity ratio. From this similarity ratio, you can then find the volume. Since 9 and 16 are squared to find the areas, take the square roots to find the similarity ratio. square root of 9 is 3 and square root of 16 is 4. Now that you have their similarity ratios, cube them both because volume is a cubed measurement. See that? Matricked was correct, it might have just seemed to be missing something to bridge the gap of understanding, if in fact there was one.

OpenStudy (anonymous):

whats the answer to that question?